OpenStudy (sundaesunday):

Help please? Since 2006, the amount of money spent at restaurants in a certain country has increased at a rate of 5% each year. In 2006​, about ​$320 billion was spent at restaurants. If the trend​ continues, about how much will be spent at restaurants in 2020?

OpenStudy (wolf1728):

Here's the formula you need: y(t) = a × e ^ (k*t) where y(t) is ending amount a=beginning amount e=2.718281828 k=.05 t=14 years (2020 minus 2006) 2020 amount = 320,000,000 * 2.71828^(k*t) 2020 amount = 320,000,000 * 2.71828^.7 2020 amount = 320,000,000 * 2.0137517593 2020 amount = 644,400,563